OpenStudy (anonymous):

Can someone help me figure this out please? A salsa recipe uses green pepper, onion, and tomato in the extended ratio 1 : 3 : 9. How many cups of onion are needed to make 117 cups of salsa? o 3 o 27 o 9 o 81

OpenStudy (anonymous):

add all the ingredients and divide it with the cups of salsa it will be the number of one quantity of any ingredient and if u want to get cups of onion multiply the ratio of onion and the one quantity of ingredient.

OpenStudy (anonymous):

you need 27 cups 1+3+9=13 127/13=9 9*3=27

OpenStudy (anonymous):

I understand. Thank you! Would you mind checking another problem I did please?

OpenStudy (anonymous):

yse

OpenStudy (anonymous):

Red and gray bricks were used to build a decorative wall. The number of red bricks over the number of gray bricks was 5/2. There were 175 bricks used in all. How many red bricks were used? o 25 o 125 o 50 o 35

OpenStudy (anonymous):

I believe it is 35.

OpenStudy (anonymous):

2+5=7 175/7=25 25*5=125

OpenStudy (anonymous):

Oh okay. Thank you very much!

OpenStudy (anonymous):

your welcome